OpenStudy (anonymous):

how bout this.... f(x)=x^2+x and g(x)=2x+1 find g^-1(x)

hero (hero):

means, find the inverse of g(x)

hero (hero):

\[g^{-1} = -\frac{1}{2x}\]

hero (hero):

Just making the form look better

OpenStudy (anonymous):

how do i simplify that

hero (hero):

Simplify? Why would you try to simplify that?

OpenStudy (anonymous):

and wat if it says g^-1(5)

hero (hero):

Then replace x with 5 and evaluate

OpenStudy (anonymous):

yes irinka just plug in the value of x if it is given on x

hero (hero):

replace x with 5, then evaluate. Don't be confused.

OpenStudy (anonymous):

so the answer would be\[g ^{-1}=\left(\begin{matrix}-1 \\ 10\end{matrix}\right)\]

OpenStudy (anonymous):

ok and (f o g)(2) would be 30?

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

what is f o g ?

OpenStudy (anonymous):

the equaion on the very top...simplify for f of g

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it means f*g (x^2+1) * (2x+1) 2x^3 + x^2 + 2x + 1 2x^3 + x^2+2x+1 is the answer

hero (hero):

irinka 17, you should clarify your question. If it is a separate question, post it as a separate question.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

11minutes ago